# Probabilistically safe controllers based on control barrier functions and scenario model predictive control

### v1

@@ Abstract
Control barrier functions (CBFs) offer an efficient framework for designing real-time safe controllers. However, CBF-based controllers can be short-sighted, resulting in poor performance, a behavior which is aggravated in uncertain conditions. This motivated research on safety filters based on model predictive control (MPC) and its stochastic variations. MPC deals with safety constraints in a direct manner, however, its computational demands grow with the prediction horizon length. We propose a safety formulation that solves a finite horizon optimization problem at each time instance like MPC, but rather than explicitly imposing constraints along the prediction horizon, we enforce probabilistic safety constraints by means of CBFs only at the first step of the horizon. The probabilistic CBF constraints is transformed in a finite number of deterministic CBF constraints via the scenario based methodology. Capitalizing on results on scenario based MPC, we provide distribution-free, a priori guarantees on the system’s closed loop expected safety violation frequency. We demonstrate our results through a case study on unmanned aerial vehicle collision free position swapping, and provide a numerical comparison with recent stochastic CBF formulations.
@@ I Introduction
Ensuring safe interactions between autonomous systems and humans in shared environments is increasingly relevant. Safe formulation in real-time can often be defined through forward set invariance, where control actions keep the system within a safe region over time. An emerging method using this concept is Control Barrier Functions [2], which has demonstrated interesting results [3], [1]. CBFs provide a principled approach to guarantee safety, while being suitable across various applications, including deterministic robotic systems settings [2], guaranteeing safety of learning methods [23], and providing safety assurances under stochastic conditions [8], which is relevant in uncertain environments.
A typical safety formulation uses quadratic programming (QP) optimization, aiming to determine least restrictive control actions. In particular, the norm of the difference between the control effort and some desired nominal controller [3], constrained by CBF inequalities is solved at each time instance. CBF usage has also been investigated in uncertain environments under the robust [21] and stochastic lenses [8].
While CBFs are computationally efficient for safety analysis, they can be short-sighted or overly conservative in control invariance calculations. To address this, predictive safety filters have emerged as a promising alternative, often used as an “add-on” to existing control strategies [18]. These approaches leverage model predictive control (MPC) to produce smoother trajectories, reduce safety filter interventions, and lower control effort [16], [4], [17]. However, these benefits come with increased computational demands, especially as the prediction horizon increase.
@@ II Related results on MPC
We will consider the design of a real-time safe controller for discrete time linear systems. At each time-step a finite horizon stochastic model predictive control (SMPC) problem will be formulated. The first component of the associated optimal input sequence will be applied to the system and the horizon will be rolled as typically performed in MPC. The first input component of each sequence will be subject to probabilistic safety constraints encoded by chance constraints. As such, the closed-loop input sequence (the concatenation of the first input components of each finite horizon problem) will be (probabilistically) safe by construction. To deal with the chance constraints present in the SMPC setting we follow a scenario based approach. To formalize this procedure we first revisit some results on stochastic and scenario based MPC.
@@ II-A Stochastic Model Predictive Control
where for each time instance $t\in\mathbb{N}$, $x_{t}\in\mathbb{R}^{n}$ denotes its state and $u_{t}\in\mathbb{R}^{p}$ the control input. The system is subject to a disturbance $d_{t}\in\mathbb{R}^{d}$ that affects the system dynamics in an affine manner. Matrices $A(d_{t}),B_{u}(d_{t})$ and $B_{d}(d_{t})$ are of appropriate dimension. For each $t$, we assume that $d_{t}\in D$, where $D$ is endowed with a $\sigma$-algebra $\mathcal{D}$. Let $\mathbb{P}$ denote a probability measure over $\mathcal{D}$ according to which $d_{t}$ is distributed, and let $\mathbb{E}$ denote the associated expected value operator.
@@ III-A1 Safety constraints
We consider state constraints encoding safety, as this is represented via control barrier functions. To this end, consider a set $S$, to represent the zero-superlevel set of a function [1] denoted as $h:\mathcal{P}\subset\mathbb{R}^{n}\to\mathbb{R}$:
@@
Ensuring system trajectories remain within $S$ guarantees safety, which is equivalent to enforcing $S$’s invariance along (1) under a given disturbance. Safety conditions can be explicitly derived using discrete-time barrier functions from [1]. We assume that $\mathbb{X}$ is defined by the following constraints:
@@
In this context, we assume that $\gamma$ lies within the interval (0, 1), though a more general approach [22] could use a class $\kappa$ functional satisfying $0<\gamma(h(x^{i}_{k|t}))\leq h(x^{i}_{k|t})$. Notice that item $(i)$ in (III-A1) omits the superscript $i$ because, at time $t$ and $k=0$, $h(x_{0|t})$ is unaffected by $d^{i}_{0|t}$ due to (1). To satisfy item $(ii)$ of (III-A1), there must exist at least one $u_{k|t}$ such that the inequality holds for all scenarios $i=1,\dots,m$ at each look-ahead step $k$. These conditions restrict the solution space, reflecting the ”price of robustness” in a distribution-free setting. When both conditions in (III-A1) are met, $h:\mathcal{P}\rightarrow\mathbb{R}$ serves as a discrete-time exponential control barrier function for the scenarios.
@@ III-B Closed loop safe set violation guarantees
Following the developments of [15], we impose the following assumptions (3).
@@
Let $\omega^{i}_{t}=\{d^{i}_{k|t}\}_{k=0}^{N-1}$ and $\omega_{t}=\{\omega^{i}_{t}\}_{i=1}^{m}$ be the collection of all finite-horizon samples. Given a current state $x_{t}(\omega_{t})$ (notice the dependence on all scenarios up to that time), denote the next-step safety violation probability as $\mathbb{P}[x_{t+1}\notin\mathbb{X}|x_{t}(\omega_{t})]$. Define $\rho$ as the support rank of the constraint $\mathbb{X}$ [15], given by $\rho=p-\text{dim}\mathcal{L}$, where $\text{dim}\mathcal{L}$ is the dimension of the largest unconstrained subspace within the decision variable space of dimension $p$. If the safety constraints are linear (as in our examples), they can be represented as $A_{t}u_{0|t}\leq b(d_{t})$, where $A_{t}$ is time-dependent. If $A_{t}\in\mathbb{R}^{z\times m}$ has a rank $r$ for all $d_{t}$ and all $t$, then $\text{dim}\mathcal{L}=p-r$. In this case $\rho\leq p$ as $0\leq\text{dim}\mathcal{L}\leq p$. Following [15, Lemma 13], we can bound this probability with the following lemma.
@@
Lemma 1 provides a bound on the next-step probability of being safe. While interesting on its own, in closed-loop operation, we seek guarantees on the safety violations of the closed-loop input sequence. Following the rationale of [15, Theorem 16], we consider the expected value of the average frequency of these violations. Let $T\in\mathbb{N}$ be the number of MPC horizon rolls, then the average number of safety violations for the closed-loop input sequence is
@@ IV-A Position swapping and collision avoidance of UAVs
Dynamics: All UAVs are represented as discrete-time double integrators; for each $j=1,\ldots,4$, each UAV’s state is given by $x_{j}=[p_{x_{j}},p_{y_{j}},v_{x_{j}},v_{y_{j}}]^{T}$, where $v_{x_{j}}$ and $v_{y_{j}}$ denote the horizontal and vertical velocity components. These dynamics are affected by an additive disturbance as in (1), $B_{d}^{j}=5\times 10^{-3}\times(-1)^{j+1}[I_{2\times 2};0_{2\times 2}]$, where $I_{2\times 2}$ and $0_{2\times 2}$ denote the $2\times 2$ identity and zero matrices. The superscript $j$ denotes the matrices/vectors for UAV $j$. We assume all disturbance scenarios are independently drawn from a normal distribution $\mathcal{N}(0,1)$, with samples also independent across the horizon and from the same disturbance type.
@@
Constraints: Safety in this setting is encoded by avoiding all pairwise collisions, enforced by the two conditions in (III-A1). A discrete time CBF to achieve this was detailed in [10]; with a slight abuse of notation, the CBF $h_{i,j}$ between agents $i$ and $j$, is $h_{i,j}=\frac{\lvert p_{x_{i}}-p_{x_{j}}\rvert}{r_{1}}+\frac{\lvert p_{y_{i}}-p_{y_{j}}\rvert}{r_{2}}-1$, where $r_{1}$ and $r_{2}$, collision radii are set to $0.25m$ and $0.5m$, and we set $\gamma=0.2$. This CBF was linearized around the states $x_{i}$, $x_{j}$ at time $t$, giving rise to constraint $A_{t}u_{0|t}\leq b(d_{t})$.
@@
Scenario based MPC and probabilistic guarantees: To solve the proposed scenario based MPC we used a prediction horizon $N=3$, for better performance while maintaining safety as explained in [10]. The matrix $A_{t}$ has a rank of $6$ (full row rank) as long as the system is safe. As such, its support rank is $\rho=6$. Theorem 1 implies that to meet the level $\epsilon=0.05$, we need at least $m=119$ scenarios.
The resulting trajectories of the scenario based MPC problem are shown in Figure 1. These trajectories achieve position swapping, and UAVs avoid collision for unseen (albeit same nature) disturbance, by being robust with respect to the considered scenarios. This is seen as the minimum CBF value for all pairwise UAVs during their fight is $h_{1,2}=0.0054$. This means UAVs 1 and 2 despite being the closest to unsafety, are still safe as $h_{1,2}\geq 0$. At the same time, by Theorem 1, the expected value of the average number of safety violations is at most $0.05$. For one of the considered UAVs, Figure 2 contrasts the deterministic trajectory with the designed (probabilistically) robust one. The associated trajectories for the other UAVs are qualitatively similar.
@@
We empirically validate Theorem 1 by setting $\gamma=0.9$ and $\epsilon=0.1$, which requires generating $m=59$ scenarios according to the theorem. Over $T=90$ time steps, we calculate the empirical expectation $\widehat{\mathbb{E}}$ for the left-hand side of (7). At each time step, we generate $m=59$ scenarios, solve the scenario program, and apply the resulting input $u^{j}_{0|t}$ to each UAV ($j=1,\ldots,4$). We compute the empirical frequency of safety violations across these time steps, considering any number of collisions at a given time as one unsafe instance to avoid double counting. This process is repeated for $M=100$ independent runs, recording the empirical frequency $f_{\ell}$ for each run. Finally, we construct the empirical expectation $\widehat{\mathbb{E}}$ as
@@ IV-B Comparative study
We compare our probabilistic safety guarantees with the approach proposed in [9] via a numerical example considered therein. Consider a one-dimension linear system: $x_{t+1}=x_{t}+2+u_{t}+\sigma d_{t}$, where $d_{t}\mathtt{\sim}\mathcal{N}(0,1)$, and define the safe set as $S_{\nu}=\{x|h(x)\geq-\nu\}$, with $h(x)=10-x^{2}$ and $\nu=0$.
@@ V Concluding remarks and future work
In this study, we introduce a sample-based approach for safe model predictive control, bridging the gap between stochastic QP-CBF and safety filters. Our method avoids the pointwise safety guarantees of myopic controllers without adding excessive constraints. It provides upper bounds on the expected frequency of chance constraint violations for systems in closed-loop by integrating a scenario-based MPC framework with control barrier functions. We evaluate this approach through two numerical studies: a UAV swapping task under external disturbances and a comparison with a state-of-the-art method offering future step exit probability bounds. These studies focus on next-step safe set exits. Promising future research includes extending the ”CBF horizon” for non-convex safe sets, ensuring recursive feasibility under mission-wide probabilistic constraints as in [20], investigating probabilistic safety under changing disturbance profiles and how higher order CBFs could impact the method proposed. Additionally, the methods compared in Section IV have interesting similarities prompting further exploration.

### v2

@@ Abstract
Control barrier functions (CBFs) offer an efficient framework for designing real-time safe controllers. However, CBF-based controllers can be short-sighted, resulting in poor performance, a behaviour which is aggravated in uncertain conditions. This motivated research on safety filters based on model predictive control (MPC) and its stochastic variant. MPC deals with safety constraints in a direct manner, however, its computational demands grow with the prediction horizon length. We propose a safety formulation that solves a finite horizon optimization problem at each time instance like MPC, but rather than explicitly imposing constraints along the prediction horizon, we enforce probabilistic safety constraints by means of CBFs only at the first step of the horizon. The probabilistic CBF constraints are transformed in a finite number of deterministic CBF constraints via the scenario based methodology. Capitalizing on results on scenario based MPC, we provide distribution-free, a priori guarantees on the system’s closed loop expected safety violation frequency. We demonstrate our results through a case study on unmanned aerial vehicle collision-free position swapping, and provide a numerical comparison with recent stochastic CBF formulations.
@@ I Introduction
Ensuring safe interactions between autonomous systems and humans in shared environments is increasingly relevant. Safety formulation in real-time can often be defined through forward set invariance, where control actions keep the system within a safe region over time. An emerging method using this concept is Control Barrier Functions [1], which has demonstrated interesting results [2], [3]. CBFs provide a principled approach to guarantee safety, while being suitable across various applications, including deterministic robotic systems settings [1], guaranteeing safety of learning methods [4], and providing safety assurances under stochastic conditions [5], which is relevant in uncertain environments.
A typical safety formulation uses Quadratic Programming (QP), aiming to determine the least restrictive control actions. In particular, the norm of the difference between the control effort and some desired nominal controller [2], constrained by CBF inequalities is solved at each time instance. CBF usage has also been investigated in uncertain environments under the robust [6] and stochastic lenses [5].
While CBFs are computationally efficient for safety analysis, they can be short-sighted or overly conservative in control invariance calculations. To address this, predictive safety filters have emerged as a promising alternative, often used as an “add-on” to existing control strategies [7]. These approaches leverage model predictive control (MPC) to produce smoother trajectories, reduce safety filter interventions, and lower control effort [8], [9], [10]. However, these benefits come with increased computational demands, especially as the prediction horizon increases.
@@ II Related results on MPC
We consider the design of a real-time safe controller for discrete time linear systems. At each time-step a finite horizon stochastic model predictive control (SMPC) problem will be formulated. The first component of the associated optimal input sequence will be applied to the system and the horizon will be rolled as typically performed in MPC. The first input component of each sequence will be subject to probabilistic safety constraints encoded by chance constraints. As such, the closed-loop input sequence (the concatenation of the first input components of each finite horizon problem) will be (probabilistically) safe by construction. To deal with the chance constraints present in the SMPC setting we follow a scenario based approach. To formalize this procedure we first revisit some results on stochastic and scenario based MPC.
@@ II-A Stochastic Model Predictive Control
where at each time instance $t\in\mathbb{N}$, $x_{t}\in\mathbb{R}^{n}$ denotes the state and $u_{t}\in\mathbb{R}^{p}$ the control input. The system is subject to a disturbance $d_{t}\in\mathbb{R}^{d}$ that affects the system dynamics in an affine manner. Matrices $A(d_{t}),B_{u}(d_{t})$ and $B_{d}(d_{t})$ are of appropriate dimension. For each $t$, we assume that $d_{t}\in D$, where $D$ is endowed with a $\sigma$-algebra $\mathcal{D}$. Let $\mathbb{P}$ denote a probability measure over $\mathcal{D}$ according to which $d_{t}$ is distributed, and let $\mathbb{E}$ denote the associated expected value operator.
@@ III-A1 Safety constraints
We consider state constraints encoding safety, as this is represented via control barrier functions. To this end, consider a set $S$ that is defined by the zero super-level set of a function [3] denoted as $h:\mathcal{P}\subset\mathbb{R}^{n}\to\mathbb{R}$:
@@
Ensuring that system trajectories remain within $S$ guarantees safety, which is equivalent to enforcing $S$’s invariance along (1) under a given disturbance. Safety conditions can be explicitly derived using discrete-time barrier functions from [3]. We assume that $\mathbb{X}$ is defined by the following constraints:
@@
In this context, we assume that $\gamma$ lies within the interval (0, 1), though a more general approach [16] could use a class $\kappa$ functional satisfying $0<\gamma(h(x^{i}_{k|t}))\leq h(x^{i}_{k|t})$. Notice that item $(i)$ in (III-A1) omits the superscript $i$ because, at time $t$ and $k=0$, $h(x_{0|t})$ is unaffected by $d^{i}_{0|t}$ due to (1). To satisfy item $(ii)$ of (III-A1), there must exist at least one $u_{k|t}$ such that the inequality holds for all scenarios $i=1,\dots,m$ at each look-ahead step $k$. These conditions restrict the solution space, reflecting the “price of robustness” in a distribution-free setting. When both conditions in (III-A1) are met, $h:\mathcal{P}\rightarrow\mathbb{R}$ serves as a discrete-time exponential control barrier function for the scenarios.
@@ III-B Closed loop safe set violation guarantees
Following the developments in [11], we impose the following assumptions on (3).
@@
Let $\omega^{i}_{t}=\{d^{i}_{k|t}\}_{k=0}^{N-1}$ and $\omega_{t}=\{\omega^{i}_{t}\}_{i=1}^{m}$ be the collection of all finite-horizon samples. Given a current state $x_{t}(\omega_{t})$ (notice the dependence on all scenarios up to that time), denote the next-step safety violation probability as $\mathbb{P}[x_{t+1}\notin\mathbb{X}|x_{t}(\omega_{t})]$. Define $\rho$ as the support rank of the constraint $\mathbb{X}$ [11], given by $\rho=p-\text{dim}\mathcal{L}$, where $\text{dim}\mathcal{L}$ is the dimension of the largest unconstrained subspace within the decision variable space of dimension $p$. If the safety constraints are linear (as in our examples), they can be represented as $A_{t}u_{0|t}\leq b(d_{t})$, where $A_{t}$ is time-dependent. If $A_{t}\in\mathbb{R}^{z\times m}$ has a rank $r$ for all $d_{t}$ and all $t$, then $\text{dim}\mathcal{L}=p-r$. In this case $\rho\leq p$ as $0\leq\text{dim}\mathcal{L}\leq p$. Following [11, Lemma 13], we can bound this probability with the following lemma.
@@
Lemma 1 provides a bound on the next-step probability of being safe. While interesting on its own, in closed-loop operation, we seek guarantees on the safety violations of the closed-loop input sequence. Following the rationale of [11, Theorem 16], we consider the expected value of the average frequency of these violations. Let $T\in\mathbb{N}$ be the number of MPC horizon rolls, then the average number of safety violations for the closed-loop input sequence is
@@ IV-A Position swapping and collision avoidance of UAVs
Dynamics: All UAVs are represented as discrete-time double integrators; for each $j=1,\ldots,4$, each UAV’s state is given by $x_{j}=[p_{x_{j}},p_{y_{j}},v_{x_{j}},v_{y_{j}}]^{T}$, where $v_{x_{j}}$ and $v_{y_{j}}$ denote the horizontal and vertical velocity components. These dynamics are affected by an additive disturbance as in (1), $B_{d}^{j}=5\times 10^{-3}\times(-1)^{j+1}[I_{2\times 2};0_{2\times 2}]$, where $I_{2\times 2}$ and $0_{2\times 2}$ denote the $2\times 2$ identity and zero matrices. The superscript $j$ denotes the matrices/vectors for UAV $j$. We assume that all disturbance scenarios are independently drawn from a normal distribution $\mathcal{N}(0,1)$, with samples also independent across the horizon and from the same disturbance type.
@@
Constraints: Safety in this setting is encoded by avoiding all pairwise collisions, enforced by the two conditions in (III-A1). A discrete time CBF to achieve this was detailed in [18]; with a slight abuse of notation, the CBF $h_{i,j}$ between agents $i$ and $j$, is $h_{i,j}=\frac{\lvert p_{x_{i}}-p_{x_{j}}\rvert}{r_{1}}+\frac{\lvert p_{y_{i}}-p_{y_{j}}\rvert}{r_{2}}-1$, where $r_{1}$ and $r_{2}$ are collision radii set to $0.25m$ and $0.5m$, and we set $\gamma=0.2$. This CBF was linearized around the states $x_{i}$, $x_{j}$ at time $t$, giving rise to constraint $A_{t}u_{0|t}\leq b(d_{t})$.
@@
Scenario based MPC and probabilistic guarantees: To solve the proposed scenario based MPC we used a prediction horizon $N=3$, for better performance while maintaining safety as explained in [18]. The matrix $A_{t}$ has rank $6$ (full row rank) as long as the system is safe. As such, its support rank is $\rho=6$. Theorem 1 implies that to meet the level $\epsilon=0.05$, we need at least $m=119$ scenarios.
The resulting trajectories of the scenario based MPC problem are shown in Figure 1. These trajectories achieve position swapping, and UAVs avoid collision for unseen (albeit of the same nature) disturbance, by being robust with respect to the considered scenarios. This is seen as the minimum CBF value for all pairwise UAVs during their fight is $h_{1,2}=0.0054$. This means UAVs 1 and 2 despite being the closest to unsafety, are still safe as $h_{1,2}\geq 0$. At the same time, by Theorem 1, the expected value of the average number of safety violations is at most $0.05$. For one of the considered UAVs, Figure 2 contrasts the deterministic trajectory with the designed (probabilistically) robust one. The associated trajectories for the other UAVs are qualitatively similar.
@@
We empirically validate Theorem 1 by setting $\gamma=0.9$ and $\epsilon=0.1$, which requires generating $m=59$ scenarios according to the theorem. Over $T=90$ time steps, we calculate the empirical expectation $\widehat{\mathbb{E}}$ for the left-hand side of (7). At each time step, we generate $m=59$ scenarios, solve the scenario program, and apply the resulting input $u^{j}_{0|t}$ to each UAV ($j=1,\ldots,4$). We compute the empirical frequency of safety violations across these time steps, considering any number of collisions at a given time as one unsafe instance to avoid double counting. This process is repeated for $M=100$ independent runs, recording the empirical frequency $f_{\ell}$ for each run. Finally, we construct the empirical expectation $\widehat{\mathbb{E}}$
@@ IV-B Comparative study
We compare our probabilistic safety guarantees with [12] using the same numerical example as in [12]. Consider a one-dimensional linear system: $x_{t+1}=x_{t}+2+u_{t}+\sigma d_{t}$, where $d_{t}\mathtt{\sim}\mathcal{N}(0,1)$, and define the safe set as $S_{\nu}=\{x|h(x)\geq-\nu\}$, with $h(x)=10-x^{2}$ and $\nu=0$.
@@ V Concluding remarks and future work
In this study, we introduce a sample-based approach for safe model predictive control, bridging the gap between stochastic QP-CBF and safety filters. Our method avoids the pointwise safety guarantees of myopic controllers without adding excessive constraints. It provides upper bounds on the expected frequency of chance constraint violations for systems in closed-loop by integrating a scenario-based MPC framework with control barrier functions. We evaluate this approach through two numerical studies: a UAV swapping task under external disturbances and a comparison with a state-of-the-art method offering future step exit probability bounds. These studies focus on next-step safe set exits. Promising future research includes extending the “CBF horizon” for non-convex safe sets, ensuring recursive feasibility under mission-wide probabilistic constraints as in [23], investigating probabilistic safety under changing disturbance profiles and how higher order CBFs could impact the method proposed. Additionally, the methods compared in Section IV have interesting similarities prompting further exploration.